19. A racing car has a uniform acceleration of 4 ms–2. What distance will it cover in 10 s after start?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Initial velocity of the car, u = 0 ms–1 Acceleration, a = 4 m s–2 Time, t = 10 s We know, distance, s = ut + (1/2)at2 Therefore, distance covered by car in 10 seconds = 0 × 10 + (1/2) × 4 × 102 = 0 + (1/2) × 4 × 10 × 10 m = (1/2) × 400 m = 200 m